A standing wave with a fundamental mode wavelength of 60 cm forms in an air column open at both ends. How long is the column for the fundamental mode?
1
Expert's answer
2020-06-04T09:57:28-0400

The condition for the resonance frequency for open at both ends tube is the following:


"f = \dfrac{nv}{2L}"

where "n" is the resonance mode (for fundamental mode "n = 1"), "v = 343 m\/s" is the speed of sound in air and "L" is the length of the air column.

Expressing "L", obtain:

"L = \dfrac{nv}{2f}"

The frequency of a standing wave can be calculated as following:


"f = v/\lambda"

where "\lambda = 0.6 m" is the wavelength.

Thus, obtain:


"L = \dfrac{nv}{2f} = \dfrac{nv}{2v/\lambda} = \dfrac{n\lambda}{2} = \dfrac{1\cdot 0.6}{2} = 0.3m"

Answer. The lenght of the tube is 0.3m.